a steamer going downstream in a river covers the distance between two towns in 20 hour coming back upstream it covers this distance in 25 hours the speed of water is 4 km/h. find the distance between the two town.
(hint : first find the speed of steamer in still water.)

Answer by ankor@dixie-net.com(22740)

a steamer going downstream in a river covers the distance between two towns in 20 hour coming back upstream it covers this distance in 25 hours the speed of water is 4 km/h. find the distance between the two town.
:
Let s = speed of the boat in still water
then
(s-4) = speed upstream
and
(s+4) = speed downstream
;
Distance both ways is equal. Make a distance equation; Distance = Time * speed
Up distance = down distance
25(s-4)= 20(s+4)
:
25s - 100 = 20s + 80
25s - 20s = +100 + 80
5s = 180
s = 180/5
s = 36 mph
:
Find the distance:
Dist = time * speed
downstream: d = 20(36+4) = 800 mi
and
Upstream: d = 25(36-4) = 800 mi also